One number is four less than 2 times a number. The sum of the numbers is 17. Find the numbers.

Answer by ikleyn(52831)

.

Let x be "a number".

Then "one number" is  2x-4.


An equation for the total is

    x + (2x-4) = 17


Simplify and find x

       3x = 17 + 4 = 21

        x          = 21/3 = 7.


ANSWER.  "A number" is 7.  "One number" is  2*7 - 4 = 10.

Solved.